In 1984, the population of the world was about 4.8 billion and the yearly growth rate was 2%. The equation giving the population P in terms of the time t (in years) is P = 4.8e0.02t. Estimate the population in 2007. (If 1984 is t = 0, then 2007 is t = 23.) When will the population be 10 billion? How many years will it take for the population to double?
Answer: The population in 2007 will be billion. The populaton will be 10 billion in the year . (Round the year off.) It will take years for the population to double. (Round up to the nearest whole number for years.)

Answer by stanbon(75887)

In 1984, the population of the world was about 4.8 billion and the yearly growth rate was 2%. The equation giving the population P in terms of the time t (in years) is P = 4.8e0.02t. Estimate the population in 2007. (If 1984 is t = 0, then 2007 is t = 23.) When will the population be 10 billion? How many years will it take for the population to double?
Answer: The population in 2007 will be billion. The populaton will be 10 billion in the year . (Round the year off.) It will take years for the population to double. (Round up to the nearest whole number for years.)
-------------------
Your equation is wrong.
The equation is P(t)=4.8(1.02)^t where t is the # of years after 1984.
If the growth rate is 2% the population
after 23 years is P(23)=4.8(1.02)^23=7.57 billion
----------
When will it reach 10 billion?
10 = 4.8(1.02)^t
2.0833333.. = 1.02^t
t= [ln2.08333...]/ln(1.02]
t=37.06 years
or 1984 +37 = 2021
-------------
When will it double; ie when will it be 9.6 billion
9.6=4.8(1.02)^t
2=(1.02)^t
t=ln2 / ln1.02 = 35 years
1984=35 = 2019
=========
Cheers,
Stan H.